# Gravitational Lensing of Euler-Heisenberg Black Hole Surrounded by Perfect Fluid Dark Matter

### v2

@@ Abstract
In this work, we study the gravitating lensing Euler-Heisenberg black hole surrounded by perfect fluid dark matter. This kind of black hole solution enables us to investigate the nontrivial interplay between the dark matter effects and nonlinear electrodynamics effects (or quantum electrodynamics effects) on charged black hole systems. The important observables in gravitational lensings are calculated and discussed in our work, including the gravitational deflection angle of light, time delay of light, precession angle of massive orbit’s bound orbits, and black hole shadow radius. The results indicates that the Euler-Heisenberg black hole with a larger perfect fluid dark matter parameter could greatly reduce the gravitational deflection angle of light, time delay of light and precession angle of massive object’s bound orbit, while the nonlinear electrodynamics effects do not have large influences on these gravitational lensing observables.
@@ 1 Introduction
In the past few decades, modern astrophysical observations reveal that most of our universe is dominated by dark matter and dark energy. Particularly, the cosmic microwave background (CMB) observations suggested that $26.8\%$ of our universe is consist of dark patter, $68.3\%$ of our universe is composed of dark energy, while only $4.9\%$ of our universe is composed of baryonic matter [2, 3, 4, 5]. Theoretically, the unknown dark matter are composed of unknown particles predicted in theories beyond the Standard Model, such as axions, sterile neutrinos and weakly interactive massive particles(WIMPs). In astronomy, the dark matter have extremely strong influences on the structure and formation of galaxy clusters, and it can also dominantly affects the star motions in galaxies and galaxy clusters, producing observational flat rotational curves in spiral galaxies [6, 7, 8, 9]. Notably, recent studies also suggested that galactic dark matter halo have non-negligible influences on the gravitational lensing and shadow images of supermassive black holes [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27].
@@
The equation of state for dark matter in galaxies (around the supermassive black hole in the galaxy center) is still a mystery. Recently, a perfect fluid dark matter (PFDM) model has attracted large interests [28, 29, 30]. Despite the simple form of PFDM model, it can reflect some non-trial effects and characters for dark matter. In recent years, a number of studies investigate the PFDM effects on many aspects of black hole physics, and effective spacetime metrics for several kinds of black holes (such as Schwarzschild, RN, Kerr black holes) surrounded by PFDM have been successfully constructed [31, 32, 33, 34, 35]. Based on these effective spacetime metrics, the gravitational lensing, black hole shadow images, quasi-normal mode, thermodynamic properties of black hole system in PFDM medium can be extensively analyzed [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52].
@@
Inspired by the simplicity and success of PFDM model in black hole physics and astronomy, we choose a novel black hole solution in PFDM medium — the Euler-Heisenberg black hole surrounded by PFDM. The Euler-Heisenberg black hole, which is obtained from the Euler-Heisenberg effective action [53, 54, 55, 56], stimulated numerous studies in past few years, for its nontrivial realization on the coupling between nonlinear electrodynamics and gravity systems, especially with the quantum loop-corrections in quantum electrodynamics (QED) [57]. With these desired properties, this kind of black hole is often used to explore the nonlinear electrodynamics in black hole physics, evoking a number of studies on thermodynamic transitions, thermodynamic topology, black hole shadow, gravitational lensing, quasi-normal modes for black hole systems with nonlinear electrodynamics in recent years [58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 84, 80, 81, 82, 83, 89, 86, 87, 91, 88, 90, 85]. The nonlinear electrodynamics effects (or QED effects) on gravitation are extremely important issues in high energy scale, and they may have potential influences on the evolution of our universe (especially at the beginning stage), such as the inflationary theory in Big Bang, the dynamic evolution of dark matter and dark energy, and the cosmological accelerating expansion of universe. Therefore, it is interesting to study the interplay between dark matter and nonlinear electrodynamics in gravitational systems. The Euler-Heisenberg black hole surrounded by PFDM could provide us a simply way to investigate their influences on charged black holes. Recently, several analytical solution for Euler-Heisenberg black hole in the presence of dark matter medium have been constructed [92, 93, 94, 95], which enable us to give a comprehensive exploration on the interplay dark matter effects and nonlinear electrodynamics effects in black hole systems.
@@
In this work, we carry out an investigation on the gravitational lensing of Eular-Heisenberg black hole surrounded by PFDM. The gravitational lensing observables provide an avenue to reveal the features of gravitational field and intrinsic properties of black holes through high precision observational data [96, 97]. An exploration of gravitational lensing observables for Eular-Heisenberg black hole surround by PFDM could enrich our knowledge on the gravitational field for charged black hole systems, especially with the influences coming from the dark matter effects and nonlinear electrodynamics effects (or QED effects). Concretely, in the present work, we calculate the gravitational deflection angle of light, time delay of light, precession angle of massive object’s bound orbits for Euler-Heisenberg black hole surrounded by PFDM. There are significantly important observables in the gravitational lensing observations, which have been frequently used to test various aspects of black hole physics and gravitational theories in both strong gravitational field limit and weak gravitational field limit. Furthermore, stimulated by the rapid development of black hole shadow researches since the capture of black hole images by Event Horizon Telescope (EHT) [98, 99, 100], the black hole critical shadow radius for Euler-Heisenberg black hole in the presence of PFDM is also discussed in our work.
@@
This rest of the present work is organized in the following way. Section 2 briefly introduces the spacetime metric generated by Euler-Heisenberg black hole surrounded by PFDM. In section 3, the theoretical approach employed in calculating the gravitational lensing observables is described in details. In section 4, numerical results and discussions on the gravitational lensing observables (gravitational deflection angle of light, time delay of light, trajectory and precession angle of massive object’s bound orbit, black hole shadow critical radius) for Euler-Heisenberg black holes surrounded by PFDM are presented. Summary and perspectives are given in section 5. Throughout this work, the natural unit $G=c=1$ is adopted.
@@ 2 Euler-Heisenberg Black Hole Surrounded by Perfect Fluid Dark Matter
The Euler-Heisenberg black hole becomes attractive in recent year for its ability to include the nonlinear electrodynamics effects (or QED effects). This kind of black hole solution can be solved from the gravitational field equation and the Euler-Heisenberg effective action [53], and it is one of the most important realization on the coupling between nonlinear electrodynamics and charged black hole systems.
@@
Recently, several analytical solution for Euler-Heisenberg black hole in the presence of dark matter medium have been constructed [93, 92, 94, 95]. Particularly, S.-J. Ma et al. derived the effective spacetime metric for an Euler-Heisenberg black hole surrounded by PFDM, through solving the gravitational field equation coupled with nonlinear electromagnetic field (with QED effects included) and dark matter field (assumed to be PFDM) [92]. The Lagrangian of a gravitational system coupled with nonlinear electromagnetic field and dark matter field can be written as [92, 94]
$S=\frac{1}{4\pi}\int\sqrt{-g}d^{4}x\cdot\bigg{[}\frac{R}{4}+\mathcal{L}_{\text{EM}}+\mathcal{L}_{\text{PFDM}}\bigg{]}$ (1)
and the nonlinear electromagnetic field’s Lagrangian density is given by
$\mathcal{L}_{\text{EM}}=-F+\frac{a}{2}F^{2}+\frac{7a}{8}G^{2}$ (2)
@@
where $F=\frac{1}{4}F_{\mu\nu}F^{\mu\nu}$ is electromagnetic strength tensor and $G=\frac{1}{4}F_{\mu\nu}*F^{\mu\nu}$ (the $*F_{\mu\nu}$ is the dual form of the electromagnetic strength tensor $F_{\mu\nu}$). The above Lagrangian leads to the gravitational field equation
$R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}=8\pi T_{\mu\nu}$ (3)
@@
with the energy-momentum tensor contributed from nonlinear electromagnetic field and dark matter field. To derive a spherically symmetric analytical solution for black hole, the energy-momentum tensors for PFDM and nonlinear electromagnetic field are restricted to the following explicit form in the spherical coordinate [92]
$T^{\mu}_{\nu}$ $=$ $(T^{t}_{t},\ T^{r}_{r},\ T^{\theta}_{\theta},\ T^{\phi}_{\phi})$ (4)
$=$ $\frac{1}{4\pi}\bigg{(}-\frac{Q^{2}}{2r^{4}}+\frac{aQ^{4}}{8r^{8}},\ -\frac{Q^{2}}{2r^{4}}+\frac{aQ^{4}}{8r^{8}},\ \frac{Q^{2}}{2r^{4}}-\frac{3aQ^{4}}{8r^{8}},\ -\frac{Q^{2}}{2r^{4}}+\frac{aQ^{4}}{8r^{8}}\bigg{)}\ \ \text{Electromagnetic Field}$
@@
$=$ $\bigg{(}\frac{\lambda_{\text{DM}}}{8\pi r^{3}},\ \frac{\lambda_{\text{DM}}}{8\pi r^{3}},\ -\frac{\lambda_{\text{DM}}}{16\pi r^{3}},\ -\frac{\lambda_{\text{DM}}}{16\pi r^{3}}\bigg{)}\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \text{Dark Matter Field}$
@@
where $M$ is the black hole mass, $Q$ is the black hole electric charge, $a$ is a parameter which measures the strength of nonlinear electrodynamics effects (or QED effects). The parameter $\lambda_{\text{DM}}$ denotes the influences of PFDM, which is proportional to the dark matter density. Eventually, by solving the gravitational field equation in the spherical coordinates, the spacetime metric for an Euler-Heisenberg black hole surrounded by PFDM is
$ds^{2}=-f(r)dt^{2}+\frac{1}{f(r)}dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2})$ (5)
@@
$f(r)=1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}-\frac{aQ^{4}}{20r^{6}}+\frac{\lambda_{\text{DM}}}{r}\cdot\ln\frac{r}{|\lambda_{\text{DM}}|}$ (6)
@@
In the absence of dark matter (namely $\lambda_{\text{DM}}=0$), this metric represents an Euler-Heisenberg black hole with nonlinear electrodynamics effects. On the other hand, when $a=0$, the nonlinear electrodynamics effects vanish and the spacetime gives a charged Reissner-Nordström (RN) black hole surrounded by PFDM. In the cases of $\lambda_{\text{DM}}=0$ and $a=0$, the spacetime recovers to a simple RN black hole spacetime.
@@
In the presence of PFDM and nonlinear electrodynamics effects, the horizon of charged black hole can be greatly influenced. It is commonly known that the classical charged black hole spacetime (such as RN spacetime) could produce naked singularities when electric charge exceeds its maximal allowed value. However, the Euler-Heisenberg black hole with a nonzero nonlinear electrodynamics parameter $a>0$ and a non-vanishing PFDM parameter $\lambda_{\text{DM}}\neq 0$ would no longer produce naked singularities for arbitrary black hole electric charge (see Appendix A). Therefore, in the rest of this work, it is safely to change black hole parameters in the calculation of gravitational lensing observables for Euler-Heisenberg black hole surrounded by PFDM, without worrying about the annoying naked singularities. The horizons of Euler-Heisenberg black hole surrounded by PFDM are discussed in the Appendix A.
@@
As an theoretical investigation, we only consider the dark matter field and nonlinear electromagnetic field in this work. The influences from other complex matter fields (such as disk structure in galaxies composed of luminous stars, dusts and plasma in galaxies) are not included in our work. Since the energy-momentum tensor of the dark matter fluid is spherically symmetric, they can form a spherically symmetric halo structure around the central supermassive black hole. In this work, we assume that the event horizon of Euler-Heisenberg black hole (surrounded by PFDM) $r_{H}$, the scale of dark matter halo $r_{\text{halo}}$, and the impact parameter for particle orbits $b$ satisfy
$r_{H}\ll b\ll r_{\text{halo}}$ (7)
@@
Furthermore, it should be mentioned that the PFDM model which gives rise to the black hole solution in (6) is a very simplified perfect fluid model. In such simplified model, the energy-momentum tensor of PFDM contains only one free parameter $\lambda_{\text{DM}}$, so the energy density $\rho$ and pressure $P$ of the isotropic dark matter fluid are directly linked with $\lambda_{\text{DM}}$ via expression (4). The spacetime metric of black holes in this simple dark matter model satisfies $g_{tt}\times g_{rr}=-1$. Recently, some studies have explored more complex perfect fluid models for dark matter. In these dark matter fluid models, the energy-momentum tensor of dark matter fluid contains the energy density $\rho$ and the pressures in radial direction and tangential direction (parameterized by $P_{r}$ and $P_{t}$), and they are connected through the equation of state for dark matter fluids [101, 102, 103, 104]. Especially, some researchers are interested in the Tolman-Oppenheimer-Volkoff (TOV) equation of state for dark matter fluids [104, 105]. The charged black hole systems solved form these dark matter models are more complex, and the $g_{tt}\times g_{rr}=-1$ relation for spacetime metric is no longer valid. The studies of more complex charged black hole solutions from dark matter model is beyond the scope of our present work.
3 Theoretical Methods to Obtain Gravitational Lensing Observables

### v3

@@ Abstract
In this work, we study the gravitational lensing of Euler-Heisenberg black hole surrounded by perfect fluid dark matter. This kind of black hole solution enables us to investigate the nontrivial interplay between the dark matter effects and nonlinear electrodynamics effects (or quantum electrodynamics effects) on charged black hole systems. The important observables in gravitational lensings are calculated and discussed in our work, including the gravitational deflection angle of light and time delay of light. Additionally, we also explore the massive orbit’s bound orbits (and their precession angles) and black hole shadow radius for Euler-Heisenberg black hole in the presence of dark matter. The results indicate that the Euler-Heisenberg black hole with a larger perfect fluid dark matter parameter could greatly reduce the gravitational deflection angle of light, time delay of light and precession angle of massive object’s bound orbit, while the nonlinear electrodynamics effects do not have large influences on these observables.
@@ 1 Introduction
In the past few decades, modern astrophysical observations reveal that most of our universe is dominated by dark matter and dark energy. Particularly, the cosmic microwave background (CMB) observations suggested that $26.8\%$ of our universe is consist of dark matter, $68.3\%$ of our universe is composed of dark energy, while only $4.9\%$ of our universe is made up of baryonic matter [2, 3, 4, 5]. Theoretically, the dark matter is composed of unknown particles predicted in theories beyond the Standard Model, such as axions, sterile neutrinos and Weakly Interacting Massive Particles (WIMPs). In astronomy, the dark matter has extremely strong influences on the structure and formation of galaxy clusters, and it can also dominantly affect the star motions in galaxies and galaxy clusters, producing observational flat rotational curves in spiral galaxies [6, 7, 8, 9]. Notably, recent studies also suggested that galactic dark matter halos have non-negligible influences on the gravitational lensing and shadow images of supermassive black holes [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27].
@@
The equation of state for dark matter in galaxies (around the supermassive black hole in the galaxy center) is still a mystery. Recently, a perfect fluid dark matter (PFDM) model has attracted great interest [28, 29, 30]. Despite the simple form of PFDM model, it can reflect some non-trivial effects and properties of dark matter. In recent years, a number of studies investigate the PFDM effects on many aspects of black hole physics, and effective spacetime metrics for several kinds of black holes (such as Schwarzschild, RN, Kerr black holes) surrounded by PFDM have been successfully constructed [31, 32, 33, 34, 35]. Based on these effective spacetime metrics, the gravitational lensing, black hole shadow images, quasi-normal mode, thermodynamic properties of black hole system in PFDM medium can be extensively analyzed [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54].
@@
Inspired by the simplicity and success of PFDM model in black hole physics and astronomy, we choose a novel black hole solution in PFDM medium — the Euler-Heisenberg black hole surrounded by PFDM. The Euler-Heisenberg black hole, which is obtained from the Euler-Heisenberg effective action [56, 55, 57, 58, 59], stimulated numerous studies in past few years, for its nontrivial realization on the coupling between nonlinear electrodynamics and gravity systems, especially with the quantum loop-corrections in quantum electrodynamics (QED) [60]. With these desired properties, this kind of black hole is often used to explore the nonlinear electrodynamics in black hole physics, evoking a number of studies on thermodynamic transitions, thermodynamic topology, black hole shadow, gravitational lensing, quasi-normal modes for black hole systems with nonlinear electrodynamics in recent years [61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 87, 83, 84, 85, 86, 93, 90, 91, 95, 92, 94, 88, 89]. The nonlinear electrodynamics effects (or QED effects) on gravitation are extremely important issues in high energy scale, and they may have potential influences on the evolution of our universe (especially at the beginning stage), such as the inflationary theory in Big Bang, the dynamic evolution of dark matter and dark energy, and the cosmological accelerating expansion of universe. Therefore, it is interesting to study the interplay between dark matter and nonlinear electrodynamics in gravitational systems. The Euler-Heisenberg black hole surrounded by PFDM could provide us a simple way to investigate its influences on charged black holes. Recently, several analytical solutions for Euler-Heisenberg black hole in the presence of dark matter medium have been constructed [96, 98, 97, 99], which enable us to give a comprehensive exploration on the interplay between dark matter effects and nonlinear electrodynamics effects in black hole systems.
@@
In this work, we carry out an investigation on the gravitational lensing of Euler-Heisenberg black hole surrounded by PFDM. The gravitational lensing, which is caused by spacetime curvature and could result in distortion or deflection effects on photon propagation, offers us an avenue to reveal the features of gravitational field and intrinsic properties of black holes through high precision observational data [100]. The Einstein’s general relativity, alternative gravity theories and astrophysical models can be tested from gravitational lensings [100, 101, 102]. In recent years, significant progress has been made on exploring gravitational lensing in large numbers of systems [103, 104, 105, 106, 107, 109, 110, 108]. Particularly, the dark matter in galaxies and dark energy in our universe can be studied through gravitational lensing observations [111, 112, 113]. The amounts of dark matter and the properties of dark matter halos in galaxies and galaxies clusters can be strongly constrained by gravitational lensing observations [114]. Furthermore, recent developments showed that the electrodynamics and plasma medium in galaxies may greatly influence the gravitational lensing [115, 116, 117, 118]. Therefore, an exploration of gravitational lensing observables for Euler-Heisenberg black hole surrounded by PFDM could enrich our knowledge on the gravitational field for charged black hole systems, especially with the influences coming from the dark matter effects and nonlinear electrodynamics effects (or QED effects). Concretely, in the present work, we calculate the gravitational deflection angle of light and time delay of light for Euler-Heisenberg black hole surrounded by PFDM. They are significantly important observables in gravitational lensing, which have been frequently used to test various aspects of black hole physics and gravitational theories in both strong gravitational field limit and weak gravitational field limit.
In addition to gravitational lensing observables, other observational quantities could also enrich our knowledge on black hole systems and gravitational fields. The precession of massive object’s bound orbits near a massive celestial body is extremely important in understanding the central gravitational sources [119, 120, 121, 122]. For example, the anomalous precession of Mercury’s orbit is a renowned example in testing the Einstein’s general relativity [123, 124], and the motion of S stars around Sgr A* provides a pathway to study the property of supermassive black hole in our galaxy [125, 126, 127]. In our work, the massive object’s bound orbits and their precession angles for Euler-Heisenberg black hole surrounded by PFDM are calculated and analyzed. Furthermore, the black hole shadow and black hole’s optical images provide direct ways to probe the various effects (eg., strong gravitational field effects, electromagnetic effects, matter fields’ effects) on black hole systems. A number of researches on black hole shadow have been stimulated since the capture of black hole images by Event Horizon Telescope (EHT) [128, 129, 130, 131]. In the presented work, the black hole critical shadow radius for Euler-Heisenberg black hole in the presence of PFDM is also discussed.
@@
The rest of the present work is organized in the following way. Section 2 briefly introduces the spacetime metric generated by Euler-Heisenberg black hole surrounded by PFDM. In section 3, the theoretical approach employed in calculating the gravitational lensing observables is described in details. In section 4, numerical results and discussions on the gravitational lensing observables (gravitational deflection angle of light and time delay of light), trajectory and precession angle of massive object’s bound orbit, black hole shadow radius for Euler-Heisenberg black holes surrounded by PFDM are presented. Summary and perspectives are given in section 5. Throughout this work, the natural unit $G=c=1$ is adopted.
@@ 2 Euler-Heisenberg Black Hole Surrounded by Perfect Fluid Dark Matter
The Euler-Heisenberg black hole becomes attractive in recent years, for its ability to include the nonlinear electrodynamics effects (or QED effects). This kind of black hole solution can be solved from the gravitational field equation and the Euler-Heisenberg effective action [56], and it is one of the most important realization on the coupling between nonlinear electrodynamics and charged black hole systems.
@@
Recently, several analytical solutions for Euler-Heisenberg black hole in the presence of dark matter medium have been constructed [98, 96, 97, 99]. Particularly, S.-J. Ma et al. derived the effective spacetime metric for an Euler-Heisenberg black hole surrounded by PFDM, through solving the gravitational field equation coupled with nonlinear electromagnetic field (with QED effects included) and dark matter field (assumed to be PFDM) [96]. The action of a gravitational system coupled with nonlinear electromagnetic field and dark matter field can be written as [96, 97]
$S=\frac{1}{4\pi}\int\sqrt{-g}\cdot d^{4}x\cdot\bigg{[}\frac{R}{4}+\mathcal{L}_{\text{EM}}+\mathcal{L}_{\text{PFDM}}\bigg{]},$ (1)
and the nonlinear electromagnetic field’s Lagrangian density is given by [60]
$\mathcal{L}_{\text{EM}}=-F+\frac{a}{2}F^{2}+\frac{7a}{8}G^{2},$ (2)
@@
where $F=\frac{1}{4}F_{\mu\nu}F^{\mu\nu}$ is the conventional electromagnetic invariant in Maxwell theory and $G=\frac{1}{4}F_{\mu\nu}*F^{\mu\nu}$ (the $*F_{\mu\nu}$ is the dual form of the electromagnetic strength tensor $F_{\mu\nu}$). The above Lagrangian leads to the gravitational field equation
$R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}=8\pi T_{\mu\nu},$ (3)
@@
with the energy-momentum tensor contributed from nonlinear electromagnetic field and dark matter field. To derive a spherically symmetric analytical solution for black hole in this theory, the energy-momentum tensor of matter fields must be assumed to be a spherically symmetric distribution. In the work from S.-J. Ma et al., the energy-momentum tensor of matter fields has the following explicit form in spherical coordinates. The total energy-momentum tensor is composed of two parts: the energy-momentum of dark matter field and the energy-momentum of nonlinear electromagnetic field [96]
$T^{\mu}_{\nu}$ $=$ $(T_{\text{EM}})_{\nu}^{\mu}+(T_{\text{DM}})_{\nu}^{\mu}$
$=$ $\text{diag}\{T^{t}_{t},\ T^{r}_{r},\ T^{\theta}_{\theta},\ T^{\phi}_{\phi}\},$
$(T_{\text{EM}})_{\nu}^{\mu}$ $=$ $\frac{1}{4\pi}\cdot\text{diag}\bigg{\{}-\frac{Q^{2}}{2r^{4}}+\frac{aQ^{4}}{8r^{8}},\ -\frac{Q^{2}}{2r^{4}}+\frac{aQ^{4}}{8r^{8}},\ \frac{Q^{2}}{2r^{4}}-\frac{3aQ^{4}}{8r^{8}},\ \frac{Q^{2}}{2r^{4}}-\frac{3aQ^{4}}{8r^{8}}\bigg{\}},\ \ \text{Electromagnetic Field}$
@@
$(T_{\text{DM}})_{\nu}^{\mu}$ $=$ $\text{diag}\bigg{\{}\frac{\lambda_{\text{DM}}}{8\pi r^{3}},\ \frac{\lambda_{\text{DM}}}{8\pi r^{3}},\ -\frac{\lambda_{\text{DM}}}{16\pi r^{3}},\ -\frac{\lambda_{\text{DM}}}{16\pi r^{3}}\bigg{\}},\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \text{Dark Matter Field}$ (4)
@@
where $M$ is the black hole mass, $Q$ is the black hole electric charge, $a$ is a parameter which measures the strength of nonlinear electrodynamics effects (or QED effects). The parameter $\lambda_{\text{DM}}$ denotes the influence from PFDM, which is proportional to the dark matter density. Eventually, by solving the gravitational field equation in the spherical coordinates, the spacetime metric for an Euler-Heisenberg black hole surrounded by PFDM is
$ds^{2}=-f(r)dt^{2}+\frac{1}{f(r)}dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2}),$ (5)
@@
$f(r)=1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}-\frac{aQ^{4}}{20r^{6}}+\frac{\lambda_{\text{DM}}}{r}\cdot\ln\frac{r}{|\lambda_{\text{DM}}|}.$ (6)
@@
The motivation and detailed derivation of the metric component function $f(r)$ are reviewed in Appendix A. In the absence of dark matter (namely $\lambda_{\text{DM}}=0$), this metric represents an Euler-Heisenberg black hole with nonlinear electrodynamics effects. On the other hand, when $a=0$, the nonlinear electrodynamics effects vanish and the spacetime gives a charged Reissner-Nordström (RN) black hole surrounded by PFDM. In the cases of $\lambda_{\text{DM}}=0$ and $a=0$, the spacetime recovers to a simple RN black hole spacetime.
@@
In the presence of PFDM and nonlinear electrodynamics effects, the horizon of charged black hole can be greatly influenced. It is commonly known that the classical charged black hole spacetime (such as RN spacetime) could produce naked singularities when electric charge exceeds its maximal allowed value. However, the Euler-Heisenberg black hole with a nonzero nonlinear electrodynamics parameter $a>0$ and a non-vanishing PFDM parameter $\lambda_{\text{DM}}\neq 0$ would no longer produce naked singularities for arbitrary black hole electric charge (see Appendix B). Therefore, in the rest of this work, it is safely to change black hole parameters in the calculation of gravitational lensing observables for Euler-Heisenberg black hole surrounded by PFDM, without worrying about the annoying naked singularities. The horizons of Euler-Heisenberg black hole surrounded by PFDM are discussed in Appendix B.
@@
As a theoretical investigation, we only consider the dark matter field and nonlinear electromagnetic field in this work. The influences from other complex matter fields (such as disk structure in galaxies composed of luminous stars, dust and plasma in galaxies) are not included in our work. Since the energy-momentum tensor of the dark matter fluid is spherically symmetric, it can form a spherically symmetric halo structure around the central supermassive black hole. In this work, we assume that the event horizon of Euler-Heisenberg black hole (surrounded by PFDM) $r_{H}$, the scale of dark matter halo $r_{\text{halo}}$, and the impact parameter for particle orbits $b$ satisfy
$r_{H}\ll b\ll r_{\text{halo}}.$ (7)
@@
Furthermore, it should be mentioned that the PFDM model which gives rise to the black hole solution in (6) is a very simplified perfect fluid model. In such simplified model, the energy-momentum tensor of PFDM contains only one free parameter $\lambda_{\text{DM}}$, so the energy density $\rho$ and pressure $P$ of the isotropic dark matter fluid are directly linked with $\lambda_{\text{DM}}$ via expression (4). The spacetime metric of black holes in this simple dark matter model satisfies $g_{tt}\times g_{rr}=-1$. Recently, some studies have explored more complex perfect fluid models for dark matter. In these dark matter fluid models, the energy-momentum tensor of dark matter fluid contains the energy density $\rho$ and the pressures in radial direction and tangential direction (parameterized by $P_{r}$ and $P_{t}$), and they are connected through the equation of state for dark matter fluids [132, 20, 133, 134]. Especially, some researchers are interested in the Tolman-Oppenheimer-Volkoff (TOV) equation of state for dark matter fluids [134, 135]. The charged black hole systems solved from these dark matter models are more complex, and the $g_{tt}\times g_{rr}=-1$ relation for spacetime metric is no longer valid. The study of more complex charged black hole solutions from dark matter model is beyond the scope of our present work.
3 Theoretical Method